5. Let C be the curve in space given parametrically, by the equations. = - 31+5, y = ( -2) and 3 = + - where 0 < < I and F
Answer #1

전 3 45키 d= (2t-3)가 지 dy H3.13tyd | E-2) Le tttta dr = utest-it) ::& fido = fadatz dytydz = t-st+5) (-3)dt +ttttttt ㅔ ,432 = l